OpenStudy (anonymous):

Is it true that for a quadratic equation ax^2+bx+c, the roots of the equation will always equal -b/a?

OpenStudy (amistre64):

nope

OpenStudy (amistre64):

-b/a aint even an option...unless you get real lucky lol

myininaya (myininaya):

it will always be x=frac{-b pm sqrt{b^2-4ac}{2a} a does not equal zero

myininaya (myininaya):

that didn't work

OpenStudy (anonymous):

There is a sort of cheat way to factor it though. I'm told it's not technically mathematically sound, but I like it and it works. Multiply a and b

OpenStudy (amistre64):

lol....its the french version

OpenStudy (amistre64):

that only works for integer roots

OpenStudy (anonymous):

Then no wonder it's not mathematically sound. Ahaha

OpenStudy (amistre64):

well, for roots that dont incvolve radicals

OpenStudy (anonymous):

And no, not really. You just gotta tweak it.

OpenStudy (anonymous):

so it will equal -b/a, assuming I'm only working with integers?

myininaya (myininaya):

\[x=\frac\]

OpenStudy (amistre64):

no.....-b/a aint even on the board

OpenStudy (anonymous):

but i keep trying to find a counterexample and cant

OpenStudy (amistre64):

-b/2a is a possibility; but that aint always a root unless the graph just touches once

OpenStudy (amistre64):

if its a complete square than -b/2a ia the only root

OpenStudy (anonymous):

-b/2a is to find the vertex...

OpenStudy (anonymous):

i just want to find a counterexample to -b/a being the sum of the roots

OpenStudy (amistre64):

-b/2a is the axis of symmetry; but its also half way between roots :)

OpenStudy (anonymous):

my bad, i forgot to mention sum

OpenStudy (amistre64):

if its a complete square than -b/2a ia the only root

OpenStudy (amistre64):

that was weird lol

myininaya (myininaya):

so you want to find a counterexample: how about x^2-5x+6=0 -b/a=5 but there are two solutions and neither of them are 5 the solutions are x=3 and x=2

OpenStudy (anonymous):

3+2=5, forgot to mention that the sum of the roots = -b/a

myininaya (myininaya):

oh ok

OpenStudy (amistre64):

the quad formula says; stand in the middle at -b/2a....now look left and right an equal distance of sqrt(...)/2a

myininaya (myininaya):

i have a proof

OpenStudy (anonymous):

?

OpenStudy (anonymous):

lol, did you just do that?

myininaya (myininaya):

yes does it make sense?

OpenStudy (anonymous):

yea, do you know how it can be useful though?

myininaya (myininaya):

you can use it to sort of check yourself instead of plugging your answers back into the equation to check you can just see they add up to be -b/a thanks for pointing this out i never realized it before

OpenStudy (anonymous):

lol, your welcome XD

myininaya (myininaya):

the square roots cancel because one was postive and one was negative just like 5+(-5)=0 so you completely understand? thats amazing well unless you are passed algebra by now

OpenStudy (anonymous):

I am. My sis just brought this to my attention and I never realized it before. XD

myininaya (myininaya):

its really cool :)

myininaya (myininaya):

by way it works for all numbers